OpenStudy (anonymous):

tell me how to graph this on a coordinate plane: y is greater than or equal to 2/3x-2

OpenStudy (anonymous):

\[y \ge \frac{ 2 }{ 3}x-2\]

jimthompson5910 (jim_thompson5910):

Now you're graphing the line y = 2/3x - 2, but this time it's a solid line and you're shading above the line.

OpenStudy (anonymous):

so if it greater than or equal to or less than or equal to,it means, a solid line. if it is greater you shade above the line... right?

jimthompson5910 (jim_thompson5910):

yes, you are absolutely correct

jimthompson5910 (jim_thompson5910):

if there's an "or equal to" added in, the line is solid because you're including the boundary

OpenStudy (anonymous):

oh i think i am catching on now a bit

jimthompson5910 (jim_thompson5910):

otherwise, it's dashed to mean you're excluding the boundary

OpenStudy (anonymous):

okay, now they through in a little curve ball. the next one is: y> -8

jimthompson5910 (jim_thompson5910):

y > -8 has the boundary line y = -8 and this boundary line is a dashed line

jimthompson5910 (jim_thompson5910):

So graph this dashed line, then shade above the dashed line

OpenStudy (anonymous):

can you show me the graph, just to make sure mine is correct?

jimthompson5910 (jim_thompson5910):

Sure thing, one moment

OpenStudy (anonymous):

okay thank you!

jimthompson5910 (jim_thompson5910):

see attached

OpenStudy (anonymous):

last but not least: 2x + 3y < 9 which is y< -2/3x +3 .. so i graph this line: y= -2/3x +3 with a dashed line and shade below it. right?

jimthompson5910 (jim_thompson5910):

btw, I'm using geogebra to make all these graphs Geogebra: http://www.geogebra.org/cms/

jimthompson5910 (jim_thompson5910):

you are 100% correct, nice work

OpenStudy (anonymous):

okay well thank you for all your help c: you are such a life saver. someone else tried to explain it but it was just too technical. thanks so much

jimthompson5910 (jim_thompson5910):

you're welcome